# Post-hoc Stochastic Concept Bottleneck Models

### v2

@@ Appendix B Comparing Different Intervention Strategies and Policies
In Section˜2, we decomposed a concept intervention into the selection of concepts for intervention and updating their values. The former part we denote as intervention policy and the latter - as intervention strategy. In the following we describe and evaluate different possible choices. We consider two intervention policies: random concept selection (as in koh2020concept) and concept uncertainty policy, which selects the concept whose predicted probability is closest to 0.5 (that is, the most uncertain concept). For intervention strategies we consider four options:
@@
    Empirical Percentile Strategy sets the value of the intervened-on concept to the 5th (if it is 0) or 95th (if it is 1) percentile of the empirical distribution of concept predictions made by the model. koh2020concept use it for intervening in CBM with soft concept encoding.
@@
    Hard Strategy introduced by havasi2022addressing sets concept probabilities to 0 or 1. In the case of PSCBM, where concepts are represented by their logits, one cannot directly use the logits of 0-1 probabilities, as they are infinite. Instead, we set concepts to the logits of $\varepsilon$ for absent concepts and $1-\varepsilon$ for present concepts, where $\varepsilon$ is a small positive number.
@@
    where $\mathcal{S}$ is the subset of concepts that have been selected for intervention, $\bm{\mu}_{\mathcal{S}}$ and $\bm{\Sigma}_{\mathcal{S},\mathcal{S}}$ are the predicted concept mean and covariance of this concept subset, $c_{i}$ are the concept values selected by the user, $d$ is the dimensionality of the distribution, i.e. the number of concepts, and $\alpha$ is the requested confidence of the confidence region. In this problem we seek to find concept logits $\eta^{\prime}_{\mathcal{S}}$ which maximize the log-probability of intervened concept values, and stay within the $\alpha$-confidence region of the predicted distribution, that is, the model’s prediction shouldn’t be completely disregarded (first constraint). The log-probability of this distribution, after being multiplied by -2, asymptotically follows the $\chi^{2}$ distribution [silvey1975statistical], which gives rise to the first inequality. The other two inequalities follow from the requirement that the new concept logits shouldn’t move away from the ground truth: if $c_{i}=1$, $\eta_{i}$ should increase, and if $c_{i}=0$, it should decrease. This last strategy can only be used in the context of SCBM, while the three others can be used by both SCBM and CBM.
@@ Appendix C Pseudocode for Interventions Training
Algorithm 1 Loss calculation for one iteration of training with random interventions
@@ Appendix D Results with amortized covariance
In this appendix, we show additional results for the studied models using an amortized covariance matrix instead of a globally learnt one as in the main manuscript. The amortized covariance, as introduced in SCBM Vandenhirtz2024stochastic, makes the $\bm{\mu}$ and $\bm{\Sigma}$ learnt per instance, i.e. amortized by the input $\bm{x}$}. We show results for CBM and SCBM as baselines and for a PSCBM trained without interventions. In Table˜3 we show test-time results for target and concept accuracy, as well as target and concept accuracy AUC for interventions with probability uncertainty policy. In Figure˜3, we show the respective intervention curves. We observe that without interventions, the performance of PSCBM is better than thus of the baselines. With interventions, it scores the highest on target accuracy AUC, and on par with the rest on concept accuracy AUC. Intervention curves confirm these results. However, in the curve for target accuracy, one can observe that after the first few interventions, the SCBM improved faster than the PSCBM. This is important because in a real-world scenario it is expected that only a few concept interventions would be made.
@@ Appendix E Implementation Details
All models have been implemented in PyTorch [ansel2024pytorch] and optimized using Adam [kingma2015adam]. To apply weight decay, we used the optimizer AdamW, which decouples weight decay from gradient computation [loshchilov2019decoupledweightdecayregularization]. All models use a ResNet-18 encoder. The remaninig parts including the concept mean and covariances, and the target predictor are all linear layers. All baselines have been trained with three random initializations. All PSCBM versions have been evaluated with each of the three CBMs with three random seeds, which amounts to a total number of nine evaluations. The computations have been done on a cluster composed mostly of NVIDIA GeForce RTX 2080 GPUs, and every job used a single GPU and two CPUs. When we train PSCBM with interventions, we use 20 random masks per datapoint per epoch, and found that a larger number does not lead to better performance. We intervene on 20% of the total number of concepts, and their values at training time are set according to the Hard Strategy, introduced in Appendix˜B. We performed hyperparameter optimization with respect to learning rate scheduling, initial learning rate, and weight decay factor. In Table˜4 we report the optimal values based on validation loss after the last training epoch. It should be noted that PSCBMa is not very sensitive to weight decay and that for PSCBMg the advantage of step learning rate over cosine schedule was minimal, albeit statistically significant. Similarly, for CBM and for both SCBM variants, all four tested hyperparameter combinations achieved similar performance.
Table 4: Optimal hyperparameter values for each model

### v4

@@ Appendix B Comparing Different Intervention Strategies and Policies
In Section˜2, we decomposed a concept intervention into the selection of concepts for intervention and updating their values. The former part we denote as intervention policy and the latter - as intervention strategy. In the following we describe and evaluate different possible choices. We consider two intervention policies: random concept selection (as in Koh et al. (2020)) and concept uncertainty policy, which selects the concept whose predicted probability is closest to 0.5 (that is, the most uncertain concept). For intervention strategies we consider four options:
@@
    Empirical Percentile Strategy sets the value of the intervened-on concept to the 5th (if it is 0) or 95th (if it is 1) percentile of the empirical distribution of concept predictions made by the model. Koh et al. (2020) use it for intervening in CBM with soft concept encoding.
@@
    Hard Strategy introduced by Havasi et al. (2022) sets concept probabilities to 0 or 1. In the case of PSCBM, where concepts are represented by their logits, one cannot directly use the logits of 0-1 probabilities, as they are infinite. Instead, we set concepts to the logits of $\varepsilon$ for absent concepts and $1-\varepsilon$ for present concepts, where $\varepsilon$ is a small positive number.
@@
    where $\mathcal{S}$ is the subset of concepts that have been selected for intervention, $\bm{\mu}_{\mathcal{S}}$ and $\bm{\Sigma}_{\mathcal{S},\mathcal{S}}$ are the predicted concept mean and covariance of this concept subset, $c_{i}$ are the concept values selected by the user, $d$ is the dimensionality of the distribution, i.e. the number of concepts, and $\alpha$ is the requested confidence of the confidence region. In this problem we seek to find concept logits $\eta^{\prime}_{\mathcal{S}}$ which maximize the log-probability of intervened concept values, and stay within the $\alpha$-confidence region of the predicted distribution, that is, the model’s prediction shouldn’t be completely disregarded (first constraint). The log-probability of this distribution, after being multiplied by -2, asymptotically follows the $\chi^{2}$ distribution (Silvey, 1975), which gives rise to the first inequality. The other two inequalities follow from the requirement that the new concept logits shouldn’t move away from the ground truth: if $c_{i}=1$, $\eta_{i}$ should increase, and if $c_{i}=0$, it should decrease. This last strategy can only be used in the context of SCBM, while the three others can be used by both SCBM and CBM.
@@ Appendix C Pseudocode for Interventions Training
Algorithm 1 Loss calculation for one iteration of training with random interventions.
@@ Appendix D Results with amortized covariance
In this appendix, we show additional results for the studied models using an amortized covariance matrix instead of a globally learnt one as in the main manuscript. The amortized covariance, as introduced in SCBM Vandenhirtz et al. (2024), makes the $\bm{\mu}$ and $\bm{\Sigma}$ learnt per instance, i.e. amortized by the input $\bm{x}$}. We show results for CBM and SCBM as baselines and for a PSCBM trained without interventions. In Table˜3 we show test-time results for target and concept accuracy, as well as target and concept accuracy AUC for interventions with probability uncertainty policy. In Figure˜3, we show the respective intervention curves. We observe that without interventions, the performance of PSCBM is better than thus of the baselines. With interventions, it scores the highest on target accuracy AUC, and on par with the rest on concept accuracy AUC. Intervention curves confirm these results. However, in the curve for target accuracy, one can observe that after the first few interventions, the SCBM improved faster than the PSCBM. This is important because in a real-world scenario it is expected that only a few concept interventions would be made.
@@ Appendix E Implementation Details
All models have been implemented in PyTorch (Ansel et al., 2024) and optimized using Adam (Kingma and Ba, 2015). To apply weight decay, we used the optimizer AdamW, which decouples weight decay from gradient computation (Loshchilov and Hutter, 2019). All models use a ResNet-18 encoder. The remaninig parts including the concept mean and covariances, and the target predictor are all linear layers. All baselines have been trained with three random initializations. All PSCBM versions have been evaluated with each of the three CBMs with three random seeds, which amounts to a total number of nine evaluations. The computations have been done on a cluster composed mostly of NVIDIA GeForce RTX 2080 GPUs, and every job used a single GPU and two CPUs. When we train PSCBM with interventions, we use 20 random masks per datapoint per epoch, and found that a larger number does not lead to better performance. We intervene on 20% of the total number of concepts, and their values at training time are set according to the Hard Strategy, introduced in Appendix˜B. We performed hyperparameter optimization with respect to learning rate scheduling, initial learning rate, and weight decay factor. In Table˜4 we report the optimal values based on validation loss after the last training epoch. It should be noted that PSCBMa is not very sensitive to weight decay and that for PSCBMg the advantage of step learning rate over cosine schedule was minimal, albeit statistically significant. Similarly, for CBM and for both SCBM variants, all four tested hyperparameter combinations achieved similar performance.
Table 4: Optimal hyperparameter values for each model.